OpenStudy (anonymous):

Need help will give medel!!! What is the interquartile range of the data? 100, 120, 130, 188, 196, 220, 265, 300 A. 200 B. 117.5 C. 242.5 D. 145

OpenStudy (anonymous):

242.5

OpenStudy (anonymous):

umm idk what the interquartile range is but the range is the largest number minus the lower number so 200 would be the range but idk if that's right because idk what interquartile range is

OpenStudy (anonymous):

mine is correct im 100% positive

OpenStudy (texaschic101):

100,120,130,188,196,220,265,300 median (Q2) = (188 + 196)/2 = 384/2 = 192 Q1 = (120 + 130)/2 = 250/2 = 125 Q3 = (220 + 265)/2 = 485/2 = 242.50 Interquartile range is Q3 - Q1 242.50 - 125 = 117.50 (B)

OpenStudy (anonymous):

but only cause (265 + 220)/2 = 242.5

OpenStudy (texaschic101):

answer is B

OpenStudy (anonymous):

nope its A because if u know math like I do then yea I have all As